June 3, 2015 Calibration
GlueX Calibration Meeting
Wednesday, June 3, 2015
11:00 am, EDT
JLab: CEBAF Center, F326
You can connect using BlueJeans using the meeting number 630 804 895 . (Click "Expand" to the right for more details -->):
- Make sure you have created a BlueJeans account via your JLab CUE account using this link:
- http://jlab.bluejeans.com (You should only need to do this once)
- Meeting ID: 630804895
- (you may need to type this in, depending how you connect)
- If connecting via Web Browser: click this link (no passcode is needed):
- If connecting via iOS or Android App:
- Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting
- If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"
- Dial: 1 888 240 2560 (US or Canada only)
- or, List of International Numbers
- If connecting via Polycom unit:
- Dial 220.127.116.11 or bjn.vc
- Enter meeting ID above
- Use *4 to unmute
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2015/ .
- Subdetector Reports
- Calibration Processing
Attending: Sean, Amiran (NU); Simon, Curtis, Mark I., Luke, Nathan (JLab); Naomi (CMU)
We had a small meeting due to the ongoing JLab User's Meeting and HUGS.
- Items that were discussed: Prospects for running in the Fall; effectiveness of our rho photoproduction slides (shown ~5-6 times at the Users Meeting); importance of looking at the data and measuring what other particles we can find
- Mike S. has tested his Millipede CDC alignment procedure using MC, and has started looking at real data. Another focus is understanding the differences between Simon's results (which have some straws with very large offsets) and Mike's (which do not). Both use different track fitting and hit selection procedures.
- There was some additional discussion on finding problem channels in the CDC.
- Other to-do items include updating CDC and FDC overall gain scales and updating the simulation code.
- Beni is still developing his TOF ADC calibrations. He has also given some updated timing constants to Simon, which are giving better results - Simon sees pion and proton bands in the ST/TOF beta v. p distribution.
- Eric (via email)
- The propogation time calibrations are well understood, and attention is now shifted to the attenuation length calibrations. These calibrations will then need to be iterated before adding them to the CCDB.
- Tagger calibration work has slown down, pending Alex Somov's analysis of the PS magnet map, which will greatly help these efforts.
- He has also been looking into prospects for measuring the total flux in the spring. The prospects do not look good, but he has developed a list of what needs to be done for the various methods to work. This discussion is proceeding in detail in the photon beam working group.
- Naomi - Sean and David will discuss adding the CDC drift time monitoring to the online tomorrow.
- Mark, Sean, and Justin have iterated on the conditions for the simulations, and Mark is now running simulation production.
- There are some long running jobs that are about to time out. It's assumed that they are due to our pushing of pythia beyond its comfort zone (i.e. E(gamma) < 3 GeV), but the log files should give a more definite answer.
- Mark showed his system for archiving and displaying information about production jobs on the queue.
- Calibration Processing
- Sean is developing the calibration processing suite using the new JLab SWIF toolset. The plan is to compare its functionality with that of Mark's toolset, and give regular feedback and feature suggestions to the JLab IT team.
- Sean is also updating the offline monitoring infrastructure and working with Kei to track down problems with the Spring data offline monitoring jobs.